ELECTRONIC SPECTROSCOPY OF Li(NH_{3})_{4}

\underline{N. BHALLA}, L. VARRIALE, N. M. TONGE, A. M. ELLIS, Department of Chemistry, University of Leicester, Leicester, LE1 7RH, United Kingdom.

Li(NH_{3})_{4} has been proposed as a key entity in lithium-ammonia solutions, but its spectral signature has so far proved impossible to distinguish from other species in these solutions. Here we present the first electronic spectrum of Li(NH_{3})_{4} in the gas phase in an effort to understand the electronic structure of this important cluster. This spectrum was recorded using mass-selective depletion spectroscopy. Strong absorption is observed in the near-infrared and the band system is assigned to the Ã^{2}T_{2} - X̃^{2}A_{1} transition in a nominally tetrahedral complex. However, an extensive vibrational progression in a low frequency bending vibration is indicative of a substantial Jahn-Teller effect in the excited electronic state. Our observations confirm a recent theoretical prediction that the electronic spectrum of Li(NH_{3})_{4} will strongly overlap with the spectrum of the solvated electron in liquid ammonia.